OpenStudy (mackenzie2013):

Find the volume of a sphere with a diameter of 17.2 cm, both in terms of π and as a number. Use 3.14 for π. Round your answers to the nearest tenth.

OpenStudy (anonymous):

v=4/3pi*r3

OpenStudy (mackenzie2013):

17.2 times pi?

OpenStudy (zale101):

How did you get that?

OpenStudy (zale101):

Volume of a sphere is \(\Large \frac{4}{3}\pi r^3\) r is the radius

OpenStudy (zale101):

Your question has the diameter. Step 1: Change diameter to radius. Do you remember how to change diameter to radius? Step 2: Plug the radius into the volume formula 4/3 *pi* r^3